OpenStudy (anonymous):

(5t-6)(5t+6) multiply and simplify

OpenStudy (anonymous):

(5t-6)(5t+6) = (5t)^2- (6)^2 = 25t^2-36 ans

OpenStudy (anonymous):

25t^2-36=0

OpenStudy (anonymous):

expand using (a+b)(a-b)=a^2-b^2